### v1

@@ Methods
where $\Delta f_{n{\bf k}}=f_{n{\bf k}}(t)-f_{n{\bf k}}^{(0)}$ is the change of the occupation function with $f_{n{\bf k}}^{(0)}$ the (equilibrium) Fermi-Dirac distribution before excitation. The time-dependent electron distribution function is obtained within the time-dependent Boltzmann equations (TDBE) [69, 70] as implemented in the EPW code. The TDBE is solved using an excited-state Ansatz for the electrons described by a high-temperature Fermi-Dirac distribution $f_{n{\bf k}}(t=0)=[{\rm exp}((\varepsilon_{n{\bf k}}-\mu)/k_{\rm B}T_{\rm exc})+1]^{-1}$, where the excitation temperature $T_{\rm exc}$ is set to match the energy transferred to the system by an incident pump with fluence $F$, as discussed in Ref. [38]. The TDBE are solved through second-order Runge-Kutta time-stepping with a time interval of 1 fs, within an energy window of 4 eV around the Fermi energy. The coherent phonon amplitude obtained from Eqs. (22) and (23) is illustrated in Fig. S3 (a) (single-pulse excitation).
Based on this framework, the maximum coherent-phonon amplitude $Q_{\nu}^{\rm max}(F)$ is estimated by considering the static limit of Eq. 22 ($\partial_{t}Q^{\nu}=0$) leading to $Q_{\nu}^{\rm max}(F)=D_{\nu}/\omega^{2}_{\nu}$. This value constitutes an upper theoretical limit to the maximum lattice displacement that the crystal can undergo for a given electronic excitation $\Delta f_{n\mathbf{k}}$.
To simulate the coherent lattice dynamics induced by two distinct laser pulses, the coherent phonon equation of motion is solved by considering the action of two driving forces:
$D^{\nu}_{\rm dp}(t;\Delta t_{p_{1}-p_{2}})=D_{\nu}(t)+D_{\nu}(t-\Delta t_{p_{1}-p_{2}})\theta(\Delta t_{p_{1}-p_{2}})\quad,$ (24)
@@
where $\Delta t_{p_{1}-p_{2}}$ is the time delay between the two pump pulses. The coherent phonon amplitudes obtained from the time propagation of Eqs. ((22)-(24)) is illustrated in Figs. S3 (b-d) for different values of the time delay $\Delta t_{p_{1}-p_{2}}$ between the two pulses. The frequency renormalization proportional to the coherent phonon amplitude observed in experiments can be phenomenologically accounted for by substituting the bare phonon frequency $\omega_{\nu}$ with the dressed frequency $\Omega_{\nu}=\omega_{\nu}-\beta|Q_{\nu}|$ in Eq. (22), where $\beta$ is a positive fitting parameter.
@@
  • Acknowledgements. This work was funded by an ETH Postdoctoral Fellowship and a Swiss National Science Foundation (SNSF) Postdoctoral Fellowship (TMPFP2 217303). The authors would like to thank M. Bauer, A. Lindenberg, S. L. Johnson, D. Soranzio, M. Savoini, and K. S. Burch for helpful discussions as well as J. Hecht and S. Reitz for technical support.
  • Author contributions. The project was conceived by J.G.H. Experiments and data analysis were conducted by J.G.H., with contributions from A.C. and T.L. Q.M. carried out the symmetry analysis and related DFT calculations. C.E. and F.C. performed the DFT calculations on the energy-dependent electron-phonon coupling. The manuscript was written by J.G.H. M.F. provided overall guidance and support. All authors discussed the results and commented on the manuscript.
@@
Figure 1: $\mid$ Ultrafast structural dynamics in ferroelectric Td-WTe2. a, Schematic of the experimental setup. b, Structural configurations of WTe2 and lattice modes connecting the Td and 1TT’ polymorphs to the paraelectric high-symmetry phase. Grey background structures denote the atomic configuration of the high-symmetry paraelectric phase, included for comparison. c, Illustration of structural dynamics on the excited-state potential energy surface following photo-excitation. d, Polarization-dependent SHG intensity at selected pump–probe delays $\Delta t_{\text{p-pr}}$. The incident polarization of the fundamental is varied, with the analyzer fixed in the horizontal orientation. e, Normalized, delay-dependent SHG efficiency (polarizer: vertical; analyzer: horizontal) as a function of absorbed laser fluence $F$. Traces are vertically offset for clarity. f, Sliding mode amplitude $A_{\text{s}}$ and lifetime $\tau_{\text{s}}$ as a function of absorbed fluence $F$. g, Sliding mode frequency and quasi-static suppression of SHG intensity as functions of absorbed fluence $F$.
@@
Figure 2: $\mid$ Ab initio modeling of sliding amplitude saturation. a, State-resolved electron-phonon coupling strength at the sliding mode frequency at two electronic temperatures $T_{\text{elec}}=2000\,\text{K}$ and $T_{\text{elec}}=4000\,\text{K}$. Red (blue) indicates positive (negative) contributions to sliding mode excitation with more saturated colors corresponding to stronger coupling. Along the energy axis, the coupling strength is weighted by the change in the Fermi–Dirac distribution $\Delta f_{\text{elec}}$ at the respective temperature. b, Calculated maximum amplitudes $Q_{\text{max}}$ of the sliding mode and two higher-frequency modes of Td-WTe2 as a function of absorbed fluence $F$. Triangular data points represent experimentally determined sliding mode amplitudes for comparison.
@@
Figure 3: $\mid$ Coherent control and amplification of sliding motion. a, Schematic of amplitude control of displacively excited coherent phonons using double-pulse excitation. b, Normalized SHG intensity as a function of pump–probe delay $\Delta t_{\text{p-pr}}$ and double-pulse delay $\Delta t_{\text{p1-p2}}$ for a combined absorbed fluence of $2.1\,\text{mJ}\,\text{cm}^{-2}$. Red, yellow, and violet arrows indicate the double-pulse delays corresponding to the pump–probe traces shown in c. c, Pump–probe SHG traces at selected double-pulse delays, demonstrating coherent control of the sliding mode amplitude. d, Simulated sliding mode amplitude $Q$ as a function of $\Delta t_{\text{p-pr}}$ and $\Delta t_{\text{p1-p2}}$, assuming an amplitude-dependent sliding frequency.e, Comparison of pump–probe SHG traces for $\Delta t_{\text{p1-p2}}=0\,\text{ps}$ and $\Delta t_{\text{p1-p2}}=8.3\,\text{ps}$. f, Sliding mode amplitude after the second optical excitation as a function of $\Delta t_{\text{p1-p2}}$. Dashed black line, expected behavior in coherent control experiments without amplitude saturation (see Methods Section); red line, fitted model incorporating amplitude suppression at short double-pulse delays; violet line, exponential component of the model used to extract the timescale of transient sliding mode suppression. g, Comparison of vibrational amplitudes extracted from pump–probe ($P$) and pump–pump–probe (PPP) measurements. Black line, estimated vibrational amplitude assuming a linear dependence of $A_{\text{s}}$ on the excited carrier density.
@@
Figure 4: $\mid$ High-amplitude vibrational spectroscopy in targeted excited states. a, Sliding mode amplitude $A_{\text{s}}$, damping $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$, and frequency $f_{\text{s}}$ as a function of the double-pulse delay $\Delta t_{\text{p1-p2}}$ for a combined absorbed fluence of $2.1\,\text{mJ}\,\text{cm}^{-2}$. Light-colored shades, $1\sigma$-confidence intervals of the fits used to extract parameters from the two-dimensional dataset shown in Fig. 3b. b, Sliding mode damping constant $\gamma_{\text{s}}$ as a function of the sliding mode amplitude $A_{\text{s}}$ for two different combined fluences in double-pulse experiments. Black line and markers, values obtained in pump-probe experiments (see Fig. 1f); solid red and violet lines, linear models fitted to the data to determine $(\Delta\gamma/\Delta A)_{\text{s}}$; errorbars, $1\sigma$-confidence intervals of the fits used to extract $A_{\text{s}}$ and $\gamma_{\text{s}}$. c, Low-energy phonon band structure of WTe2.
Figure 5: $\mid$ Modulation of nonlinear optical properties by high-amplitude coherent sliding motion. a, Fourier-transform of the two-dimensional dataset in Fig. 3b along the $\Delta t_{\text{p-pr}}$ axis. The data has been normalized to the maximum spectral density at the fundamental sliding mode frequency. b, Fourier spectra extracted from a at three selected double-pulse delays. c, The $2\omega_{\text{s}}$-contribution to the SHG signal as a function of the double-pulse delay $\Delta t_{\text{p1-p2}}$. The fundamental sliding mode amplitude determined from fits (see Figs. 3 and 4) is shown for comparison. d, Sliding-mode-amplitude-dependent $2\omega_{\text{s}}$-contribution.
@@
Fig. S1 $\mid$ Polarization-dependent SHG signal in double-pulse and single-pulse experiments a, Comparison of SHG signals from WTe2 as a function of the delay between two pump pulses for parallel (violet) and cross-polarized (red) configurations (combined absorbed fluence $F=2.1\,\text{mJ}\,\text{cm}^{-2}$). The probe pulse interrogates the sample at a fixed delay of $\Delta t_{\text{p-pr}}=100\,\text{ps}$, where coherent phonon oscillations are negligible. For parallel-polarized pump pulses, interference at the sample surface leads to a delay-dependent modulation of energy absorption, resulting in a corresponding variation of the SHG signal. In contrast, interference effects are strongly suppressed for cross-polarized pump pulses. Consequently, all double-pulse excitation experiments were performed using cross-polarized pump pulses. b, SHG intensity measured at $\Delta t_{\text{p-pr}}=100\,\text{ps}$ for two different polarizer–analyzer configurations, shown as a function of the polarization angle of a single pump pulse (absorbed fluence $F=1.6\,\text{mJ}\,\text{cm}^{-2}$; $0^{\circ}:P\parallel a$-axis; $90^{\circ}:P\parallel b$-axis of the crystal). Only a weak anisotropy is observed, indicating that in double-pulse experiments both pump pulses, despite their differing polarizations, contribute similarly to the electronic excitation of the material.
@@
Fig. S2 $\mid$ Time-dependent polarimetry measurements a–d, SHG intensity at selected pump–probe delays $\Delta t_{\text{p-pr}}$ as a function of the incident polarization angle of the fundamental light. The analyzer was fixed in the horizontal direction. Violet circles, experimental data; violet lines, fits based on the material’s symmetry (see Methods Section for details). e, Comparison of polarization-resolved SHG intensity just before and immediately after time zero. A pronounced change is observed in the $\chi_{zyy}$ component (at $90^{\circ}$), while the $\chi_{\text{mix}}$ component (at $0^{\circ}$) remains largely unaffected. f, Delay-dependent evolution of the SHG tensor components $\chi_{zyy}$ (violet) and $\chi_{\mathrm{mix}}$ (red), extracted from fits to the time-resolved polarimetry data. Upon photoexcitation, $\chi_{zyy}$ exhibits a rapid drop to a quasi-static level with superimposed oscillations at the sliding-mode frequency. In contrast, $\chi_{\mathrm{mix}}$ shows neither a fast initial suppression nor a comparable long-lived change.
@@
Fig. S3 $\mid$ Time-resolved simulations of the coherent phonon excitation. a Time-dependent coherent phonon amplitude $Q_{\nu}$ for a single driving pulse, obtained from the solution of Eqs. ((22)-(23)), considering a single driving force marked in brown. b-c Coherent phonon amplitude and driving force in the presence of two-pulse excitation as described via Eqs. ((22)-(24)) for different time delays $\Delta t_{p_{1}-p_{2}}$ between the two pump pulses. Panel b illustrates the case in which the time delay coincides with one full phonon period, resulting in resonant excitation. In panels c and d, the time delay is set to 1.25 and 1.5 phonon periods. In these cases, the second pulse results in the destructive interference of the coherent phonon, as revealed by the lower coherent phonon amplitude compared to the resonant case.
@@
Fig. S4 $\mid$ Double-pulse excitation of coherent phonons. Coherent phonon amplitude as a function of time delay between two pump pulses $\Delta t_{p_{1}-p_{2}}$ and of the pump-probe delay $\Delta t_{p_{2}-p_{r}}$. a Coherent phonon amplitude obtained from Eq. 22 in absence of damping ($\gamma_{\nu}=0$). b Coherent phonon amplitude obtained from Eq. 22 in presence of damping. c Coherent phonon amplitude obtained by considering the frequency renormalization according to $\Omega_{\nu}=\omega_{\nu}-\beta|Q_{\nu}|$ in Eq. (22).

### v2

@@ Methods
where $\Delta f_{n{\bf k}}=f_{n{\bf k}}(t)-f_{n{\bf k}}^{(0)}$ is the change of the occupation function with $f_{n{\bf k}}^{(0)}$ as the (equilibrium) Fermi-Dirac distribution before excitation. The time-dependent electron distribution function is obtained within the time-dependent Boltzmann equations (TDBEs) [caruso_ultrafast_2022, pan_strain-induced_2024] as implemented in the EPW code. The TDBEs are solved using an excited-state ansatz for the electrons described by a high-temperature Fermi-Dirac distribution $f_{n{\bf k}}(t=0)=[{\rm exp}((\varepsilon_{n{\bf k}}-\mu)/k_{\rm B}T_{\rm exc})+1]^{-1}$, where the excitation temperature $T_{\rm exc}$ is set to match the energy transferred to the system by an incident pump pulse with fluence $F$, as discussed in Ref. [emeis_coherent_2025]. The TDBE are solved through second-order Runge-Kutta time stepping with a time interval of 1 fs, within an energy window of 4 eV around the Fermi energy. The coherent-phonon amplitude obtained from Eqs. 16 and 17 is illustrated in Fig. S3a (single-pulse excitation). Based on this framework, the maximum coherent-phonon amplitude $Q_{\nu}^{\rm max}(F)$ is estimated by considering the static limit of Eq. 16 ($\partial_{t}Q^{\nu}=0$) leading to $Q_{\nu}^{\rm max}(F)=D_{\nu}/\omega^{2}_{\nu}$. This value constitutes an upper theoretical limit to the maximum lattice displacement that the crystal can undergo for a given electronic excitation $\Delta f_{n\mathbf{k}}$. To simulate the coherent lattice dynamics induced by two distinct laser pulses, the coherent phonon equation of motion is solved by considering the action of two driving forces:
$D^{\nu}_{\rm dp}(t;\Delta t_{\text{p}_{1}-\text{p}_{2}})=D_{\nu}(t)+D_{\nu}(t-\Delta t_{\text{p}_{1}-\text{p}_{2}})\theta(\Delta t_{\text{p}_{1}-\text{p}_{2}})\quad,$ (18)
@@
where $\Delta t_{p_{1}-p_{2}}$ is the time delay between the two pump pulses. The coherent-phonon amplitudes obtained from the time propagation of Eqs.16-18 is illustrated in Figs. S3 b-d for different values of the time delay $\Delta t_{\text{p}_{1}-\text{p}_{2}}$ between the two pulses. The frequency renormalization proportional to the coherent-phonon amplitude observed in experiments can be phenomenologically accounted for by substituting the bare phonon frequency $\omega_{\nu}$ with the dressed frequency $\Omega_{\nu}=\omega_{\nu}-\beta|Q_{\nu}|$ in Eq. refeq:osci, where $\beta$ is a positive fitting parameter.
Symmetry relationships and calculation of the free-energy landscape. In the following, we again clarify the symmetry relationships between the different structural phases of WTe2 and derive the Landau free-energy landscape depicted in Fig. 1a. As elucidated in the main text, the polar Td phase belongs to the non-centrosymmetric space group $Pmn2_{1}$, while the non-polar 1T’ phase crystallizes in the centrosymmetric $P2_{1}/m$ structure. Both can be derived from a higher-symmetry reference structure belonging to the $Pnma$ space group, which serves as their common supergroup. The relevant lattice instabilities of this $Pnma$ phase correspond to two irreducible representations: the polar interlayer sliding mode $\Gamma_{2}^{-}$ and the non-polar intralayer shear mode $\Gamma_{4}^{+}$. The $\Gamma_{2}^{-}$ distortion breaks inversion symmetry and lowers the symmetry to $Pmn2_{1}$, giving rise to the ferroelectric Td phase, while the $\Gamma_{4}^{+}$ distortion leads to monoclinic symmetry ($P2_{1}/m$), corresponding to the ferroelastic 1T’ phase. We note that the interlayer sliding mode is often referred to in the literature as a ‘shear mode’. However, as evident from its irreducible representation ($\Gamma_{2}^{-}$), this mode does not break the symmetry of the elastic tensor and, therefore, is not a shear mode in the strict sense. In contrast, it is the intralayer shear mode ($\Gamma_{4}^{+}$) that is responsible for breaking an elastic symmetry.
To describe the energetics associated with these structural distortions, we construct a Landau-type free-energy expansion of the $Pnma$ reference phase in terms of the mode amplitudes $Q_{\Gamma_{2}^{-}}$ and $Q_{\Gamma_{4}^{+}}$. Up to sixth order, the expansion reads
$F=$ $F_{Pnma}+a_{\Gamma_{2}^{-}}\,Q_{\Gamma_{2}^{-}}^{2}+b_{\Gamma_{2}^{-}}\,Q_{\Gamma_{2}^{-}}^{4}+c_{\Gamma_{2}^{-}}\,Q_{\Gamma_{2}^{-}}^{6}$ (19)
$+a_{\Gamma_{4}^{+}}\,Q_{\Gamma_{4}^{+}}^{2}+b_{\Gamma_{4}^{+}}\,Q_{\Gamma_{4}^{+}}^{4}+c_{\Gamma_{4}^{+}}\,Q_{\Gamma_{4}^{+}}^{6}$
$+\gamma\,Q_{\Gamma_{2}^{-}}^{2}Q_{\Gamma_{4}^{+}}^{2}+\eta_{1}\,Q_{\Gamma_{2}^{-}}^{4}Q_{\Gamma_{4}^{+}}^{2}+\eta_{2}\,Q_{\Gamma_{2}^{-}}^{2}Q_{\Gamma_{4}^{+}}^{4}\,.$
Here, $Q_{\Gamma_{2}^{-}}$ is the amplitude of the polar interlayer sliding mode, which directly controls the out-of-plane ferroelectric polarization, while $Q_{\Gamma_{4}^{+}}$ represents the amplitude of the intralayer shear mode, associated with monoclinic distortion. The coupling coefficients $\gamma$, $\eta_{1}$, and $\eta_{2}$ describe nonlinear phonon–phonon interactions between the two order parameters and are essential for capturing the transition pathways between the $Pnma$ phase on the one hand and the Td or 1T’ phases on the other hand. The fitted parameters entering Eq. 19 are summarized in Table 1. We compute the free energy from first principles using the VASP code [kresse_ab_1993] with the DFT+D3 van der Waals functional with Becke–Johnson damping [grimme_effect_2011]. The free-energy landscape was obtained by linearly interpolating the atomic positions between the relaxed $Pnma$, Td, and 1T’ phases on a $11\times 11$ grid, allowing the lattice parameters to relax at each interpolation step.
Parameter Value [eV]
$a_{\Gamma_{2}^{-}}$ $2.00\times 10^{-2}$
$a_{\Gamma_{4}^{+}}$ $7.13\times 10^{-3}$
$b_{\Gamma_{2}^{-}}$ $1.90\times 10^{-2}$
$\gamma$ $5.90\times 10^{-2}$
$b_{\Gamma_{4}^{+}}$ $3.72\times 10^{-3}$
$c_{\Gamma_{2}^{-}}$ $1.41\times 10^{-3}$
$\eta_{1}$ $6.14\times 10^{-3}$
$\eta_{2}$ $1.53\times 10^{-3}$
$c_{\Gamma_{4}^{+}}$ $1.73\times 10^{-4}$
Table 1: Fitted coefficients of the Landau free-energy expansion (up to sixth order) in terms of the order parameters $Q_{\Gamma_{2}^{-}}$ and $Q_{\Gamma_{4}^{+}}$. Energies are given in eV.
The effective phonon potential discussed in the main text (see Fig. 1a) is obtained by expanding the total free energy around its minimum at $(Q_{\Gamma_{2}^{-}},Q_{\Gamma_{4}^{+}})=(Q_{\Gamma_{2}^{-}}^{(0)},0)$, corresponding to the equilibrium configuration of the polar Td phase. This procedure allows us to isolate the anharmonic contributions governing the coupling between the ferroelectric and ferroelastic modes, and to map out the full free energy landscape connecting the Td, $Pnma$, and 1T’ structures. In this configuration, any modulation of of the polar interlayer sliding mode $Q_{\Gamma_{2}^{-}}$ around its equilibrium value, that is $\delta Q_{\Gamma_{2}^{-}}=Q_{\Gamma_{2}^{-}}-Q_{\Gamma_{2}^{-}}^{(0)}$, transforms according to the totally symmetric $A_{1}$ irreducible representation of the $Pmn2_{1}$ space group. This means that near the Td minimum, the effective restoring potential for such fluctuations has $A_{1}$ symmetry, corresponding to the interlayer sliding mode that directly modulates the ferroelectric polarization $P_{z}$.
@@
  • Acknowledgements. This work was funded by an ETH Postdoctoral Fellowship, a Swiss National Science Foundation (SNSF) Postdoctoral Fellowship (TMPFP2 217303), and an ETH Seed Grant. The authors would like to thank M. Bauer, A. Lindenberg, S.L. Johnson, D. Soranzio, M. Savoini, and K.S. Burch for helpful discussions as well as J. Hecht, S. Reitz, and P. Moll for technical support.
  • Author contributions. The project was conceived by J.G.H. Experiments and data analysis were conducted by J.G.H., with contributions from A.C., N.W., and T.L. Q.M. carried out the symmetry analysis and related ab initio calculations. C.E. and F.C. performed the ab initio calculations on the energy-dependent electron-phonon coupling. The manuscript was written by J.G.H. M.F. provided overall guidance and support. All authors discussed the results and commented on the manuscript.
@@
Figure 1: $\mid$ Ultrafast structural dynamics in ferroelectric Td-WTe2. a, Calculated two-dimensional potential energy surface of WTe2 spanned by interlayer sliding and intralayer shear displacements. b, Structural configurations of WTe2 and lattice modes connecting the Td and 1T’ polymorphs to the paraelectric high-symmetry phase. Dark grey circles, tungsten atoms; light brown circles, tellurium atoms; grey background structures denote the atomic configuration of the high-symmetry paraelectric phase, included for comparison. c, Polarization-dependent SHG intensity at selected pump–probe delays $\Delta t_{\text{p-pr}}$. The incident polarization of the fundamental is varied, with the analyzer fixed in the horizontal orientation. d, Normalized, delay-dependent SHG efficiency (polarizer: vertical; analyzer: horizontal) as a function of absorbed laser fluence $F$. Traces are vertically offset for clarity. e, Sliding-mode amplitude $A_{\text{s}}$ and lifetime $\tau_{\text{s}}$ as a function of absorbed fluence $F$. f, Sliding-mode frequency and quasi-static suppression of SHG intensity as functions of absorbed fluence $F$.
@@
Figure 2: $\mid$ Ab initio modeling of sliding-amplitude saturation. a, State-resolved electron-phonon coupling strength $g$ at the sliding-mode frequency at two electronic temperatures $T_{\text{elec}}=2000\,\text{K}$ and $T_{\text{elec}}=4000\,\text{K}$. Red (blue) indicates positive (negative) contributions to the sliding mode excitation with more saturated colors corresponding to stronger coupling. Along the energy axis, the coupling strength $g$ is weighted by the change in the Fermi–Dirac distribution $\Delta f_{\text{elec}}$ at the respective temperature. b, Calculated maximum amplitudes $Q_{\text{max}}$ of the sliding mode and of two higher-frequency modes of Td-WTe2 as a function of absorbed fluence $F$. Triangular data points represent the sliding-mode amplitudes determined in our experiments for comparison.
@@
Figure 3: $\mid$ Coherent control and amplification of sliding motion. a, Schematic of amplitude control of displacively excited coherent phonons using double-pulse excitation. b, Normalized SHG intensity as a function of pump–probe delay $\Delta t_{\text{p-pr}}$ and double-pulse delay $\Delta t_{\text{p1-p2}}$ for a combined absorbed fluence of $2.1\,\text{mJ}\,\text{cm}^{-2}$. Red, yellow, and violet arrows indicate the double-pulse delays corresponding to the pump–probe traces shown in (c). c, Pump–probe SHG traces at selected double-pulse delays, demonstrating coherent control of the sliding-mode amplitude. d, Simulated sliding-mode amplitude $Q$ as a function of $\Delta t_{\text{p-pr}}$ and $\Delta t_{\text{p1-p2}}$, assuming an amplitude-dependent sliding frequency. e, Comparison of pump–probe SHG traces for $\Delta t_{\text{p1-p2}}=0\,\text{ps}$ and $\Delta t_{\text{p1-p2}}=8.3\,\text{ps}$. f, Sliding-mode amplitude after the second optical excitation as a function of $\Delta t_{\text{p1-p2}}$. Dashed black line, expected behaviour in coherent control experiments without amplitude saturation (see Methods); red line, fitted model incorporating amplitude suppression at short double-pulse delays; violet line, transient suppression of the sliding mode extracted from fits to the experimental data (see Methods for details). g, Comparison of vibrational amplitudes extracted from pump–probe ($P$) and pump–pump–probe (PPP) measurements. Black line, estimated vibrational amplitude assuming a linear dependence of $A_{\text{s}}$ on the excited-carrier density.
@@
Figure 4: $\mid$ High-amplitude vibrational spectroscopy in targeted excited states. a, Sliding-mode amplitude $A_{\text{s}}$, damping $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$, and frequency $f_{\text{s}}$ as a function of the double-pulse delay $\Delta t_{\text{p1-p2}}$ for a combined absorbed fluence of $2.1\,\text{mJ}\,\text{cm}^{-2}$. Light-colored shades, $1\sigma$-confidence intervals of the fits used to extract parameters from the two-dimensional dataset shown in Fig. 3b. b, Sliding-mode damping constant $\gamma_{\text{s}}$ as a function of the sliding-mode amplitude $A_{\text{s}}$ for two different combined fluences in double-pulse experiments. Solid red and violet lines, linear models fitted to the data to determine $(\Delta\gamma/\Delta A)_{\text{s}}$; errorbars, $1\sigma$-confidence intervals of the fits used to extract $A_{\text{s}}$ and $\gamma_{\text{s}}$. c, Low-energy phonon band structure of WTe2. Red arrows indicate an exemplary pathway for the Klemens-like decay of the sliding phonon into two acoustic phonons of opposite momentum.
@@
Fig. S1 $\mid$ Polarization-dependent SHG signal in double-pulse and single-pulse experiments. a, Comparison of SHG signals from WTe2 as a function of the delay between two pump pulses for parallel (violet) and cross-polarized (red) configurations (combined absorbed fluence $F=2.1\,\text{mJ}\,\text{cm}^{-2}$). The probe pulse interrogates the sample at a fixed delay of $\Delta t_{\text{p-pr}}=100\,\text{ps}$, where coherent phonon oscillations are negligible. For parallel-polarized pump pulses, interference at the sample surface leads to a delay-dependent modulation of energy absorption, resulting in a corresponding variation of the SHG signal. In contrast, interference effects are strongly suppressed for cross-polarized pump pulses. Consequently, all double-pulse excitation experiments were performed using cross-polarized pump pulses. b, SHG intensity measured at $\Delta t_{\text{p-pr}}=100\,\text{ps}$ for two different polarizer–analyzer configurations, shown as a function of the polarization angle of a single pump pulse (absorbed fluence $F=1.6\,\text{mJ}\,\text{cm}^{-2}$; $0^{\circ}:P\parallel a$-axis; $90^{\circ}:P\parallel b$-axis of the crystal). Only a weak anisotropy is observed, indicating that in double-pulse experiments both pump pulses, despite their differing polarizations, contribute similarly to the electronic excitation of the material.
@@
Fig. S2 $\mid$ Time-dependent polarimetry measurements. a–d, SHG intensity at selected pump–probe delays $\Delta t_{\text{p-pr}}$ as a function of the polarization angle of the fundamental light. The analyzer was fixed in the horizontal direction. Violet circles, experimental data; violet lines, fits based on the material’s symmetry (see Methods for details). e, Comparison of polarization-resolved SHG intensity just before and immediately after time zero. A pronounced change is observed in the $\chi_{zyy}$ component (at $90^{\circ}$), while the $\chi_{\text{mix}}$ component (at $0^{\circ}$) remains largely unaffected. f, Delay-dependent evolution of the SHG tensor components $\chi_{zyy}$ (violet) and $\chi_{\mathrm{mix}}$ (red), extracted from fits to the time-resolved polarimetry data. Upon photoexcitation, $\chi_{zyy}$ exhibits a rapid drop to a quasi-static level with superimposed oscillations at the sliding-mode frequency. In contrast, $\chi_{\mathrm{mix}}$ shows neither a fast initial suppression nor a comparable long-lived change.
@@
Fig. S3 $\mid$ Time-resolved simulations of the coherent phonon excitation. a, Time-dependent coherent phonon amplitude $Q_{\nu}$ for a single driving pulse, obtained from the solution of Eqs. 16-17, considering a single driving force marked in brown. b-d, Coherent phonon amplitude and driving force in the presence of two-pulse excitation as described via Eqs. 16-18 for different time delays $\Delta t_{\text{p}_{1}-\text{p}_{2}}$ between the two pump pulses. Panel (b) illustrates the case in which the time delay coincides with one full phonon period, resulting in resonant excitation. In panels (c) and (d), the time delay is set to 1.25 and 1.5 phonon periods. In these cases, the second pulse results in the destructive interference of the coherent phonon, as revealed by the lower coherent phonon amplitude compared to the resonant case.
@@
Fig. S4 $\mid$ Double-pulse excitation of coherent phonons. a-c, Coherent phonon amplitude as a function of time delay between two pump pulses $\Delta t_{\text{p}_{1}-\text{p}_{2}}$ and of the pump-probe delay $\Delta t_{\text{p}_{2}-\text{p}_{r}}$. a, Coherent phonon amplitude obtained from Eq. 16 in absence of damping ($\gamma_{\nu}=0$). b, Coherent phonon amplitude obtained from Eq. 16 in presence of damping. c, Coherent phonon amplitude obtained by considering the frequency renormalization according to $\Omega_{\nu}=\omega_{\nu}-\beta|Q_{\nu}|$ in Eq. 16.
Fig. S5 $\mid$ Experimental geometry. Sketch of the experimental geometry for time-resolved SHG measurements with single- and double-pulse excitation. The coordinate system refers to the a-axis and b-axis of the WTe2 crystal.
Fig. S6 $\mid$ Comparison of amplitude-dependent damping in pump-probe and pump-pump-probe experiments. b, Sliding-mode damping constant $\gamma_{\text{s}}$ as a function of the sliding-mode amplitude $A_{\text{s}}$ for two different combined fluences in double-pulse experiments (see also Fig. 4b). Solid red and violet lines, linear models fitted to the data to determine $(\Delta\gamma/\Delta A)_{\text{s}}$; errorbars, $1\sigma$-confidence intervals of the fits used to extract $A_{\text{s}}$ and $\gamma_{\text{s}}$. Black line and markers, values obtained in pump-probe experiments (see fluence-dependent sliding mode amplitude and damping in Fig. 1e). Notably, the linear relationship between the sliding-mode amplitude and its damping is evident only in the double-pulse data.